OpenStudy (anonymous):

Find the exponential function f(x)=a^x whose graph goes through the point (3,8). a= ?

OpenStudy (anonymous):

since the function has the form: y=a^x, just plug in (3,8) to find a... 8 = a^3 can you find out what a is?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

so your function is f(x)=2^x ....:)